OpenStudy (anonymous):

Can someone help me solve this integral? I have the answer, but I need to understand how to get it. The problem is simple, (integral) of cos^-1 x dx

OpenStudy (accessdenied):

So, the integral of the inverse cosine function? Did you have any ideas on how you might begin? The style of the answer might hint at it as well.

OpenStudy (anonymous):

Honestly I really need help from the start, I was sick and missed class and just got the answer basically from notes,

OpenStudy (accessdenied):

Does integration by parts sound familiar? Or this form: \( \displaystyle \int u \ dv = uv - \int v \ du \)

OpenStudy (anonymous):

can i show u what i have?

OpenStudy (accessdenied):

Sure. :)

OpenStudy (anonymous):

I = ∫cos^-1(x)dx u = cos^-1(x), then du = -1/sqrt(1-x^2) dv = dx, then v = x I = xcos^-1(x) + ∫x/sqrt(1-x^2)dx u = 1-x^2, then du = -2xdx, and xdx = -1/2du I = xcos^-1(x) -1/2 ∫du/sqrt(u) I = xcos^-1(x) -sqrt(u) + c I = xcos^-1(x) -sqrt(1-x^2) + c Is this the proper steps?

OpenStudy (accessdenied):

Yup, that looks right to me.

OpenStudy (anonymous):

Ok cool thanks lol wasn't sure

OpenStudy (accessdenied):

Yeah, no problem! :p